There is a narrow rectangular plot, reserved for a school, in Mahuli village. The length and breadth of the plot are in the ratio 11:4. At the rate ₹100 per metre it will cost the village panchayat ₹75000 to fence the plot. What are the dimensions of the plot? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the length of the rectangular plot be 11x and breadth be 4x.
Rate of fencing per metre = ₹100
Total cost of fencing = ₹75000
Perimeter of the plot = 2(l+b) = 2(11x + 4x) = 2×15x = 30x
Total amount of fencing = (30x × 100)
According to the question, (30x × 100) = 75000
⇒3000x = 75000
⇒x = 75000/3000
⇒x=25
Length of the plot = 11x = 11 × 25 = 275m
Breadth of the plot = 4 × 25 = 100m.

Thanks